-√12(4 - 2√3) I need step by step help on that problem. I don't know how to distribute with the negative square root of 12 and I would apreaciate the help

Answer by Edwin McCravy(20055)


-√12(4 - 2√3)
Simplify √12 = √4×3 = √4×√3 = 2√3
-2√3(4 - 2√3)
-8√3 + 4√3×√3
-8√3 + 4×3
-8√3 + 12
12 - 8√3
Edwin